The market value of your house is $285,000, and you have a first mortgage balance of $92,000. If a lender requires a 65% loan-to-market value ratio, how large could your home equity loan be?
Answer #1

Market value to house = $285,000

Required loan to market value = 65%

Therefore, maximum loan amount = $285000 x 65% =$185,250

Earlier mortgage balance = $92,000

Therefore maximum loan allowed on home = 185,250 - 92000 = $93250

Therefore Equity value in home = market value of home - maximum loan allowed = 285,000 - 93250 = $ 191,750